TinaaaaaA:

IF 10 Points Equal Five points How many Points is 267??

TinaaaaaA:

ii meant question. if 10 questions

boredfr:

5 is half of ten All you have to do is multiple 267 by 2 which would be 7*2=14 60*2=120 200*2=400 which added would be 14+120+400=534